# TraceFL: Interpretability-Driven Debugging in Federated Learning via Neuron Provenance

### v3

@@ IV-C Measuring Client’s Contribution
To find the end-to-end contribution, we must accumulate neuron-level provenance, $c_{n_{j}}*p_{k}*\sum_{i=1}w_{k}^{i}*z^{i}$, of a given client’s model to derive its complete contributions toward the global model’s prediction. A client’s overall contribution to the global model prediction is determined by the sum of the client’s contribution to the neurons of the global model. Specifically, if the set of activated neurons of the global model is denoted by $n_{1},n_{2},\cdots,n_{j}$, then the total contribution ($T_{k}$) of the client $k$ can be calculated using Equation 5 as follows:
@@ V Experimental Evaluations
  • Is TraceFL equally accurate on FL of different models and architectures such as LLMs and transformers?
@@
Models and Datasets. We evaluate TraceFL on state-of-the-art and commercially used CNNs, including ResNet-18 [26] and DenseNet-121 [27], as well as the two most popular transformer models, BERT [28] and GPT [29] to demonstrate the wide applicability of TraceFL. We train ResNet and DenseNet on CIFAR-10 [49] and MNIST [50]. These network-dataset combinations are widely used in practice and serve as standardized benchmarks in practice [1, 37]. We also evaluate TraceFL on real-world medical imaging datasets, including the Colon Pathology dataset [30] and Abdominal CT dataset [31, 32], to demonstrate its usability in complex real-world FL systems. The Colon Pathology dataset contains 107,180 biomedical images representing nine classes of colon pathology, while the Abdominal CT dataset contains 58,830 images of abdominal CT scans representing 11 classes. More details about these datasets can be found in  [51, 52]. For NLP tasks, we evaluate TraceFL on BERT and GPT models trained on the DBpedia [53] and Yahoo Answers [54] datasets. The DBpedia dataset contains 560,000 training samples and 70,000 testing samples, while the Yahoo Answers dataset contains 1,400,000 training samples and 60,000 testing samples, representing 14 and 10 classes, respectively.
Data Distribution Among Clients. We use Dirichlet distribution in FL to distribute non-overlapping data points among clients in each round. This is the standard FL data distribution method proven to produce real-world distribution [55, 38, 56, 37]. The parameter ($\alpha$) in Dirichlet ranges from [0, $\infty$), determining the level of Non-IID in experiments. For instance, when $\alpha$ equals 100, it replicates uniform local data distributions, while smaller $\alpha$ values increase the probability that clients possess samples from a single class [56]. A value of 0.5 is a common practice in prior work [55, 37]. We use an even stricter parameter value of 0.3 to stress test TraceFL and demonstrate its usability in more challenging cases. Nevertheless, Section V-C1 performs sensitivity analysis by varying $\alpha$ from 0.1 to 1.
Experimental Environment. To resemble real-world FL, we deploy our experiments in Flower FL [39], running on an enterprise-level cluster comprising six NVIDIA DGX A100 [57] nodes. Each node is equipped with 2048 GB of memory, at least 128 cores, and an A100 GPU with 80 GB of memory. We vary training rounds between 15 to 80 with clients ranging from 100 to 1000, thus testing TraceFL on more configurations than any related work [56, 58]. Ten randomly selected clients participate in each round, reflecting a real-world scenario where not all the clients participate in the given round [59]. We evaluate TraceFL with FedAvg [1].
@@ V-A TraceFL’s Localization Accuracy in Correct Predictions
Identifying the clients most responsible for correct prediction is a key debugging objective that helps encourage future participation of those clients to improve the overall FL accuracy. Note that TraceFL directly does not improve the FL model accuracy. Instead, it reasons about the behavior of the FL global model which an FL developer can use to improve the FL model accuracy (e.g., selecting clients which are contributing more in the FL global model predictions.) TraceFL’s neuron provenance traces predictions back to clients trained on those labels, ranking clients by their contribution. TraceFL returns a ranked list of clients in descending order of responsibility towards a prediction, where the client with the highest score is likely to be most responsible.
We evaluate TraceFL’s localization accuracy on two real-world medical imaging datasets, two standardized image datasets, and two NLP classification datasets using ResNet, DenseNet, BERT, and GPT models resulting in over 12 FL configurations spanning a total 400 FL rounds and 4000 models. We verify if the most responsible client returned by TraceFL contains the data with the label that was correctly predicted by the global model. We measure the accuracy on 10 test inputs. Figure 2 shows TraceFL’s performance in localizing responsible clients.The X-axis represents training rounds, while the Y-axis shows the FL global model’s classification accuracy and TraceFL’s localization accuracy.
@@ 1.0 NA 100
FL’s global model can exhibit unwanted behavior (e.g.,, mispredictions) due to intentional or unintentional faults in the training data of clients. Mislabelling in training data may occur due to faulty sensors, human error in labeling data, or, in some cases, adversarial attacks [1, 2, 3, 4, 5]. Finding a client responsible for such behavior is a crucial debugging goal that helps FL developers exclude such clients from participating in future rounds to improve the global model’s quality.
To evaluate TraceFL’s localization accuracy on mispredicted labels by a global model, we design the following experiments.
Similarly to prior work on fault localization in FL [33], we select one client in an FL round and flip a specific label in its training data to make it faulty. The inclusion of such clients influences the global model to make mispredictions. For instance, in the medical dataset, we flip the label ‘Cancer-associated Stroma’ to ‘Adipose’ in the Colon Pathology dataset to reflect a faulty hospital containing incorrect label data that may occur due to misdiagnosis. Using this setting from FedDebug [33] also offers fair experiment settings that compare TraceFL’s localization accuracy with FedDebug’s, which is the most recent work on fault localization in FL.
Table I shows the results. TraceFL outperforms FedDebug significantly and can operate in cross-domain tasks of image and text classification without any change in its approach. FedDebug, by construction, is limited in two ways. First, it is only designed and evaluated on image classification tasks, while TraceFL works with both image and text classification tasks. Second, it is limited to a very specific set of non-IID data distributions, whereas TraceFL performance is consistent across varying data distributions. Nevertheless, the results show that even on image classification tasks, TraceFL outperforms FedDebug in terms of localization accuracy. For instance, in Abdominal-CT with $\alpha=1$, FedDebug’s average accuracy is 44.4% while TraceFL’s accuracy is 100%.
Takeaway. TraceFL’s achieves 99.3% average localization accuracy across 18 FL settings, whereas FedDebug’s average localization accuracy is only 32% on image classification.
@@ V-C TraceFL’s Robustness
Figure 3: TraceFL performance on different data distributions.
@@ V-C2 Differential Privacy-Enabled FL
DP Noise DP Sensitivity FL Model Accuracy TraceFL Avg. Accuracy
@@
We evaluate TraceFL’s robustness when DP is enabled in FL, using standard DP settings in FL that provide optimal privacy and model accuracy, as mentioned in prior work [60]. Table IIpresents the results of this experiment, and Figure 4 shows the impact of noise on the FL training accuracy. As expected, the FL model’s accuracy decreases when the DP noise increases and vice versa. However, TraceFL maintains its performance in DP-enabled FL. As DP adds noise to the model weights, the global model’s output is still based on its neurons’ activations on the given input. Thus, TraceFL’s working principle remains intact, and it successfully traces back to the source of the prediction based on the global model’s neuron provenance. We want to emphasize that TraceFL does not recover the individual clients’ data points. It only identifies the responsible clients in ranked order. Overall, we find that TraceFL is robust against the use of differential privacy in FL where it achieves an average localization accuracy of 99% in GPT and DBpedia dataset (Figure 4 and Table II).
@@ V-D TraceFL’s Scalability
We asses the scalability of TraceFL across three different dimensions: (1) by increase the total clients, (2) by increasing the client participation, and (2) by increasing the number of rounds. First, we vary the number of clients from 200 to 1000 and measure if TraceFL can still accurately localize the responsible client. We use the state-of-the-art neural network GPT and the DBpedia dataset. Table III presents the results of the scalability experiment. We observe that TraceFL’s performance remains consistent, with an average localization accuracy of 99% across 200 to 1000 clients over a total of 75 FL training rounds. This experiment significantly exceeds the scale of experiments performed by prior work [33].
When we vary the number of participating clients per round from 20 to 50, TraceFL’s performance remains stable, achieving 100% localization accuracy across 60 FL training rounds. Prior work has shown that even at an enterprise scale, only a few clients participate in a single FL round [36, 34]. Furthermore, we evaluate the scalability of TraceFL over up to 80 rounds with 400 clients in total. Figure 5 demonstrates that TraceFL maintains consistent performance with an average localization accuracy of 98%. These results indicate that TraceFL is scalable and can handle a large number of clients and rounds without compromising its performance.
@@ V-F Threat to Validity and Limitations
There are two primary threats to the validity of the results presented. First, in our experiments, we select a randomly selected subset of clients to participate in every FL round. A different sequence of randomly selected participating clients may alter the TraceFL’s accuracy. We mitigate this threat by performing responsible client localization on every round and then reporting the average localization accuracy across all rounds. Second, the same Dirichlet distribution ($\alpha$) may provide a different distribution of the training data across clients. Even when the value $\alpha$ is the same, the localization accuracy of TraceFL may vary slightly. We mitigate this threat by averaging the localization accuracy across rounds and also measuring the localization accuracy on different datasets and models. As with any other research, TraceFL also has limitations. TraceFL is designed for classification tasks in FL. It may not be directly applicable to non-classification tasks such as text generation in GPT.
@@ VI Related Work
Debugging in Machine Learning. As the complexity of neural network models continues to increase, the need for interpretability techniques becomes more crucial and important. Interpretability techniques are used to understand the inner workings of a neural network. These techniques try to explain the decisions made by the model, and how the model makes these decisions. This is important for many reasons, including the ability to explain which input features are important to a model’s output, to understand the model’s behavior, and to identify potential biases and errors in a trained model. Several approaches, such as Integrated Gradients [41], Gradient SHAP [42], DeepLIFT [43], Saliency [44], Guided GradCAM [45], Occlusion (also called sliding window method) [46], and LIME [47], exist which evaluate the contribution of each input feature to model’s output. For instance, Integrated Gradients [41] evaluates the contribution of each input feature by calculating the integral of gradients w.r.t. input. This is done along the path from a selected baseline to the given input. Occlusion involves replacing each contiguous rectangular region with a predetermined baseline or reference point and measuring the difference in the model’s output. This approach is based on perturbations and provides a way to evaluate the importance of input features by measuring the change in the model’s output.
@@
Almost all existing provenance and interpretability approaches are inapplicable in FL, as by design, they solve an orthogonal problem — identifying the important feature in the input responsible for a prediction instead of clients. This distinction is critical because the training data or the training process are completely inaccessible in FL. Existing approaches require access to the client’s data. Furthermore, they are only designed for a single neural network but FL global model is a mixture of clients’ models participating in the given round. Operating these techniques on FL would require us to first identify a suspicious client’s mode–a problem that TraceFL solves. Even if such techniques are applied to a client’s model, the resulting feedback is not immediately actionable and constructive. TraceFL is designed to address the limitations of the existing debugging approaches and added challenges of FL such as distributed training, inaccessibility to clients, and the mixture of models.
@@
There has been recent work on ensuring accountability in FL systems. A vast majority of solutions leverage the blockchain to ensure accountability [61, 62, 63, 64, 65]. Some of these works (BlockFLow [64], BlockFLA [65]) design an FL system that uses the Ethereum blockchain to provide accountability and monetary rewards for good client behavior. However, all these systems require utilizing the blockchain and entail significant modifications to the existing FL system, presenting a barrier to adoption. TraceFL in contrast can work with any existing system without modifications.
@@ VII Conclusion
In this paper, we introduced the concept of “neuron provenance” and developed a debugging tool, TraceFL, for Federated Learning (FL). TraceFL accurately identifies the primary contributors to a global FL model’s behavior. Our evaluations show that TraceFL achieves an impressive average localization accuracy of 99%. Additionally, TraceFL also outperforms the existing fault localization technique. We provide a reusable, functional artifact of TraceFL in Flower framework to have an immediate practical impact in real-world FL deployment. In conclusion, our work represents a significant step forward in addressing the open challenges of debugging in FL.

### v4

@@ IV-C Measuring Client’s Contribution
To find the end-to-end contribution, we must accumulate neuron-level provenance, $c_{n_{j}}*p_{k}*\sum_{i=1}w_{k}^{i}*z^{i}$, of a given client’s model to derive its complete contributions toward the global model’s prediction. A client’s overall contribution to the global model prediction is determined by the sum of the client’s contribution to the neurons of the global model. Specifically, if the set of neurons of the global model is denoted by $n_{1},n_{2},\cdots,n_{j}$, then the total contribution ($T_{k}$) of the client $k$ can be calculated using Equation 5 as follows:
@@ V Experimental Evaluations
  • Is TraceFL equally accurate on FL of different models and architectures such as CNNs and transformers (GPT)?
@@
Models and Datasets. We evaluate TraceFL on state-of-the-art and commercially used CNNs, including ResNet-18 [26] and DenseNet-121 [27], as well as the two most popular transformer models, BERT [28] and GPT [29] to demonstrate the wide applicability of TraceFL. We train ResNet and DenseNet on CIFAR-10 [45] and MNIST [46]. These network-dataset combinations are widely used and serve as standardized benchmarks in practice [1, 37]. We also evaluate TraceFL on real-world medical imaging datasets, including the Colon Pathology dataset [30] and Abdominal CT dataset [31, 32], to demonstrate its usability in complex real-world FL systems. The Colon Pathology dataset contains 107,180 biomedical images representing nine classes of colon pathology, while the Abdominal CT dataset contains 58,830 images of abdominal CT scans representing 11 classes. More details about these datasets can be found in [47, 48]. For NLP tasks, we evaluate TraceFL on BERT and GPT models trained on the DBpedia and Yahoo Answers datasets [49]. The DBpedia dataset contains 560,000 training samples and 70,000 testing samples, while the Yahoo Answers dataset contains 1,400,000 training samples and 60,000 testing samples, representing 14 and 10 classes, respectively.
Data Distribution Among Clients. We use Dirichlet distribution in FL to distribute non-overlapping data points among clients in each round. This is the standard FL data distribution method proven to produce real-world distribution [50, 38, 51, 37]. The parameter ($\alpha$) in Dirichlet ranges from [0, $\infty$), determining the level of Non-IID in experiments. For instance, when $\alpha$ equals 100, it replicates uniform local data distributions, while smaller $\alpha$ values increase the probability that clients possess samples from a single class [51]. A value of 0.5 is a common practice in prior work [50, 37]. We use an even stricter parameter value of 0.3 to stress test TraceFL and demonstrate its usability in more challenging cases. Nevertheless, Section V-C1 performs sensitivity analysis by varying $\alpha$ from 0.1 to 1. These settings inherently simulate varying degrees of label overlap among clients. To explicitly manage overlapping labels, pathological data distributions can be employed [37], as shown in Figure 1. The pathological data distribution is available in TraceFL’s artifact. Furthermore, TraceFL’s artifact contains configurable data distributions among clients and allows evaluations on varying numbers of test inputs.
Experimental Environment. To resemble real-world FL, we deploy our experiments in Flower FL [39], running on an enterprise-level cluster of six NVIDIA DGX A100 [52] nodes. Each node is equipped with 2048 GB of memory, at least 128 cores, and an A100 GPU with 80 GB of memory.
We vary training rounds between 15 to 80 with clients ranging from 100 to 1000, thus testing TraceFL on more configurations than any related work [51, 53]. Ten randomly selected clients participate in each round, reflecting a real-world scenario where not all the clients participate in the given round [54]. We evaluate TraceFL with FedAvg [1].
@@ V-A TraceFL’s Localization Accuracy in Correct Predictions
Identifying the clients most responsible for correct prediction is a key debugging objective that helps encourage future participation of those clients to improve the overall FL accuracy. Note that TraceFL directly does not improve the FL model accuracy. Instead, it reasons about the behavior of the FL global model which an FL developer can use to improve the FL model accuracy (e.g., selecting clients which are contributing more in the FL global model predictions).
TraceFL’s neuron provenance traces predictions back to clients trained on those labels, ranking clients by their contribution. TraceFL returns a ranked list of clients in descending order of responsibility towards a prediction, where the client with the highest score is likely to be most responsible.
We evaluate TraceFL’s localization accuracy on two real-world medical imaging datasets, two standardized image datasets, and two NLP classification datasets using ResNet, DenseNet, BERT, and GPT models resulting in over 12 FL configurations spanning a total 400 FL rounds and 4000 models. We verify if the most responsible client returned by TraceFL contains the data with the label that was correctly predicted by the global model. We measure the accuracy on at least 10 test inputs in each round. Figure 2 shows TraceFL’s performance in localizing responsible clients. The X-axis represents training rounds, while the Y-axis shows the FL global model’s classification accuracy and TraceFL’s localization accuracy.
@@ 1.0 NA 100
FL’s global model can exhibit unwanted behavior (e.g., mispredictions) due to intentional or unintentional faults in the training data of clients. Mislabelling in training data may occur due to faulty sensors, human error in labeling data, or, in some cases, adversarial attacks [1, 2, 3, 4, 5]. Finding a client responsible for such behavior is a crucial debugging goal that helps FL developers exclude such clients from participating in future rounds to improve the global model’s quality.
To evaluate TraceFL’s localization accuracy on mispredicted labels by a global model, we design the following experiments with ten clients. Similarly to prior work on fault localization in FL, FedDebug [33], we select one client in an FL round and flip a specific label in its training data to make it faulty. The inclusion of such clients influences the global model to make mispredictions. For instance, in the medical dataset, we flip the label ‘Cancer-associated Stroma’ to ‘Adipose’ in the Colon Pathology dataset to reflect a faulty hospital containing incorrect label data that may occur due to misdiagnosis.
Table I shows the results. TraceFL outperforms FedDebug significantly and can operate in cross-domain tasks of image and text classification without any change in its approach. This is expected since FedDebug, by construction, applies to a different problem setting, i.e., debugging the model instead of the prediction, and it primarily targets a specific set of Non-IID data distributions. Even on image classification tasks, TraceFL outperforms FedDebug in terms of localization accuracy. For instance, in Abdominal-CT with $\alpha=1$, FedDebug’s average accuracy is 44.4% while TraceFL’s accuracy is 100%.
Takeaway. TraceFL achieves 99.3% average localization accuracy across 18 FL settings, whereas FedDebug’s average localization accuracy is only 32% on image classification.
@@ V-C TraceFL’s Robustness
Figure 3: TraceFL performance on different data distributions. The X-axis represents the values of Dirichlet alpha.
@@ V-C2 Differential Privacy-Enabled FL
DP Noise DP Sensitivity FL Model Accuracy % TraceFL Avg. Accuracy %
@@
We evaluate TraceFL’s robustness when DP is enabled in FL, using standard DP settings in FL that provide optimal privacy and model accuracy, as mentioned in prior work [55]. Table II presents the results of this experiment, and Figure 4 shows the impact of noise on the FL training accuracy. As expected, the FL model’s accuracy decreases when the DP noise increases and vice versa.
However, TraceFL maintains its performance in DP-enabled FL. As DP adds noise to the model weights, the global model’s output is still based on its neurons’ activations on the given input. Thus, TraceFL’s working principle remains intact, and it successfully traces back to the source of the prediction based on the global model’s neuron provenance. We want to emphasize that TraceFL does not recover the individual clients’ data points. It only identifies the responsible clients in ranked order. Overall, we find that TraceFL is robust against the use of differential privacy in FL where it achieves an average localization accuracy of 99% in GPT and DBpedia dataset (Figure 4 and Table II).
@@ V-D TraceFL’s Scalability
We assess the scalability of TraceFL across three different dimensions: (1) by increase the total clients, (2) by increasing the client participation, and (2) by increasing the number of rounds. First, we vary the number of clients from 200 to 1000 and measure if TraceFL can still accurately localize the responsible client. We use the state-of-the-art neural network GPT and the DBpedia dataset. Table III presents the results of the scalability experiment. We observe that TraceFL’s performance remains consistent, with an average localization accuracy of 99% across 200 to 1000 clients over a total of 75 FL training rounds. This experiment significantly exceeds the scale of experiments performed by prior work [33].
When we vary the number of participating clients per round from 20 to 50, TraceFL’s performance remains stable, achieving 100% localization accuracy across 60 FL training rounds. Prior work has shown that even at an enterprise scale, only a few clients participate in a single FL round [36, 34]. Furthermore, we evaluate the scalability of TraceFL over up to 80 rounds with 400 clients in total. Figure 5 demonstrates that TraceFL maintains consistent performance with an average localization accuracy of 98%. These results indicate that TraceFL is scalable and can handle numerous clients and rounds without compromising its performance.
@@ V-F Threat to Validity and Limitations
There are two primary threats to the validity of the results. First, in our experiments, we select a random subset of clients to participate in every round. A different sequence of randomly selected participating clients may alter the TraceFL’s accuracy. We mitigate this threat by performing responsible client localization on every round and then reporting the average localization accuracy across rounds. Second, the same Dirichlet distribution ($\alpha$) may provide a different distribution of the training data across clients. Even when the value $\alpha$ is the same, the localization accuracy of TraceFL may vary slightly. We mitigate this threat by averaging the localization accuracy across rounds and also measuring the localization accuracy on different datasets and models. TraceFL is designed for classification tasks in FL and may not be directly applicable to non-classification tasks such as text generation [56] and embeddings generation [57].
@@ VI Related Work
Debugging and Interpretability in Machine Learning. As the complexity of neural network models continues to increase, the need for interpretability techniques becomes more crucial and important. Interpretability techniques are used to understand the inner workings of a neural network. These techniques try to explain the decisions made by the model, and how the model makes these decisions. This is important for many reasons, including the ability to explain which input features are important to a model’s output, to understand the model’s behavior, and to identify potential biases and errors in a trained model. Several approaches, such as Integrated Gradients [58], Gradient SHAP [59], DeepLIFT [60], Saliency [61], Guided GradCAM [62], Occlusion (also called sliding window method) [63], and LIME [64], exist which evaluate the contribution of each input feature to model’s output. For instance, Integrated Gradients [58] evaluates the contribution of each input feature by calculating the integral of gradients w.r.t. input. This is done along the path from a selected baseline to the given input. Occlusion involves replacing each contiguous rectangular region with a predetermined baseline or reference point and measuring the difference in the model’s output. This approach is based on perturbations and provides a way to evaluate the importance of input features by measuring the change in the model’s output.
Existing debugging techniques [65, 66, 67, 68, 69] are designed to identify issues and enhance the performance of a single neural network in centralized ML. These methods typically require access to training data, which is prohibited in FL. For example, NPC [68] constructs a Decision Graph using training data. Furthermore, these approaches have not been evaluated on modern neural network architectures such as Transformers.
@@
Almost all existing debugging and interpretability approaches are inapplicable in FL, as by design, they solve an orthogonal problem — identifying the important feature in the input responsible for a prediction instead of clients. This distinction is critical because the training data or the training process is completely inaccessible in FL. Existing approaches require access to the client’s data. Furthermore, they are only designed for a single neural network, but the FL global model is a mixture of clients’ models participating in the given round. Operating these techniques on FL would require us to first identify a suspicious client’s mode–a problem that TraceFL solves. Even if such techniques are applied to a client’s model, the resulting feedback is not immediately actionable and constructive. TraceFL is designed to address the limitations of the existing debugging approaches and added challenges of FL, such as distributed training, inaccessibility to clients, and the mixture of models.
FedDebug [33] introduces differential testing in FL to identify faulty clients by capturing each client’s activations for a given input and localizing the client(s) whose behavior deviates from others. Building on FedDebug, a backdoor detection technique in FL is presented in [10]. Additionally, FedGT [11] aims to identify malicious clients in FL; however, it is limited to scaling up to 15-30 clients and has not been tested on advanced architectures like GPT.
Despite their contributions, these existing methods target a narrower problem under a specialized setting. FedDebug is limited to image classification tasks using CNNs, restricting its applicability to Transformer architectures. Additionally, it is designed primarily for faulty clients and IID distributions [37], as demonstrated in Table I. In contrast, TraceFL targets a broader debugging problem using a domain-agnostic and highly accurate client localization mechanism applicable to diverse neural network architectures, data types, and distributions through its novel fine-grained neuron provenance.
@@
There has been recent work on ensuring accountability in FL systems. A vast majority of solutions leverage the blockchain to ensure accountability [70, 71, 72, 73, 74]. Some of these works (BlockFLow [73], BlockFLA [74]) design an FL system that uses the Ethereum blockchain to provide accountability and monetary rewards for good client behavior. However, all these systems require utilizing the blockchain and entail significant modifications to the existing FL system, presenting a barrier to adoption. TraceFL, in contrast, can work with any existing system without modifications.
@@ VII Conclusion
We introduce the concept of neuron provenance and developed a debugging and interpretability tool, TraceFL, for FL. TraceFL accurately identifies the primary contributors to a global model’s behavior. Our evaluations show that TraceFL achieves an impressive average localization accuracy of 99%. Furthermore, TraceFL also outperforms the existing fault localization technique. We provide a reusable functional artifact of TraceFL in the Flower framework to have an immediate practical impact in real-world FL deployment, addressing the open challenges of debugging and interpretability in FL.
Acknowledgement
We thank anonymous reviewers for providing valuable and constructive feedback to help improve the quality of this work. This work was supported in part by Amazon - Virginia Tech Initiative in Efficient and Robust Machine Learning, 4-VA, and the National Science Foundation award 2106420. We also thank the Advanced Research Computing Center at Virginia Tech and the Flower FL framework for their support in building and evaluating this work.